# An Infinite Transitivity Theorem

### v1

@@ 1.2. Outline
We say a representation $\pi:C\rightarrow\mathscr{B}(H)$ of a $C^{*}$-algebra $C$ admits infinite transitivity if for every separable subspace $K$ of $H$, the map $C\rightarrow\mathscr{B}(K)$ by $c\mapsto P_{K}\pi(c)P_{K}$ is onto, where $P_{K}$ is the projection onto $K$. This is equivalent to the statement that there is always an operator in $C$ sending a given orthonormal sequence to another given bounded sequence in $H$. The first question one might have is: Does every irreducible representation of the Calkin algebra have infinite transitivity? We produce a negative answer.
@@
Finally, we find an irreducible representation of the Calkin algebra which fails infinite transitivity, granting theorem 1. Compare this situation with those of [FA10], wherein the question of whether the commutant of $\mathscr{B}(H)$ in $\mathscr{B}(H)^{\mathscr{U}}$ is trivial also depends on the ultrafilter $\mathscr{U}$. In our case, for a diagonal pure state on $\mathscr{Q}(H)$ (or equivalently $\mathscr{B}(H)$), whether the GNS-representation admits infinite transitivity depends on the corresponding ultrafilter $\mathscr{U}$. Another recent instance of this ultrafilter-dependence phenomenon was in [ZA25].
@@ 1.4. Preliminaries
where $a_{0}\in C$, $a_{ijk}\in\tilde{C}$, $n\in\mathbb{N}$. A degree-$1$ condition ofCC is an expression of the form
@@ 2.1. Quantum Filters
The beauty of maximal quantum filters is their natural bijective correspondence with pure states, i.e. $\varphi\mapsto\mathscr{F}_{\varphi}$ is a bijection from pure states to maximal quantum filters [FA19, 5.3.8]. While the map $\varphi\mapsto\mathscr{F}_{\varphi}$ is an injection, it is virtually never a bijection, from states to quantum filters. For example, $\{1\}$ is always a quantum filter, but is only of the form $\mathscr{F}_{\varphi}$ when $A=\mathbb{C}$.
@@ 2.3. Equivalences
We state these two lemmas separately from theorem 2.8, as they apply to all states, not only pure ones. Of particular interest to operator algebraists is the fact that these states have weak normality, a property first discovered for projections on the Calkin algebra by Bice in [BI13, 6.2]. That is, decreasing sequences $(a_{n})$ in $A_{+}$ have lower bounds $a\in A_{+}$ with $\varphi(a)=\lim_{n\rightarrow\infty}\varphi(a_{n})$. Also, lemma 2.7 does not require any saturation.
@@
$\|(1-e_{\gamma})a-(1-e_{\gamma})\|$ $\leq\|1-e_{\gamma}\|\|a-a^{\prime}\|+\|e_{\gamma}a^{\prime}-a^{\prime}\|$
@@
so $(1-e_{\gamma})$ satisfies $t(x)$, so it is realized by $b\in\mathscr{F}_{\varphi}$ s.t. $b\ll a$ for all $a\in Y$. Since $\mathscr{F}$ was arbitrary, we are done.
@@
For $(4)\Leftrightarrow(7)$, we apply both directions of [BI13, 4.1], so $(4)\vee(7)\Rightarrow(4)\wedge(7)\Rightarrow(3)$.
@@
Without loss of generality, by the linearity of $\varphi$, whenever $\rho_{j}=\varphi(\cdot)$, assume $z_{j}=0$.
@@
$\rho^{\prime}_{j}=\begin{cases}\|\cdot\|,&\rho_{j}=\|\cdot\|;\\ \|a\cdot a\|,&\rho_{j}=\varphi(\cdot);\end{cases}$
fix the degree-$1$ norm type
@@
and notice $(\bar{b}_{m})$ satisfies $t^{\prime}(\bar{x})$. Since $C$ is countably degree-1 saturated, $t^{\prime}(\bar{x})$ is realized. When $\rho_{j}=\varphi(\cdot)$, we have
@@ 3. Infinite Transitivity
Let $t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\in K_{j}:j\in\mathbb{N}\}$ be a satisfiable countable type consisting of conditions above. To prove this, we will perform two conversions of the above
@@
is a satisfiable generalized countable degree-1 norm type whose realizations realize $t^{\prime}(\bar{x})$, thus realize $t(\bar{x})$. Therefore, by [FA19, 15.2.10], $t(\bar{x})$ is realized. ∎
@@
Let $K_{j}$ be a separable subspace of $H_{j}$ and numerate an orthonormal basis $(\xi_{j})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\bigoplus_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{j})$. For each $j$, there exists some P-state $\varphi_{j}$ such that there is a lift $a_{j}+L_{\varphi_{j}}=\xi_{j}$. Similarly, there is some lift $b_{j}+L_{\varphi_{j}}=T\xi_{j}$. Define the type

### v2

@@ 1.2. Outline
We say a representation $\pi:C\rightarrow\mathscr{B}(H)$ of a $C^{*}$-algebra $C$ admits infinite transitivity if for every separable subspace $K$ of $H$, the map $C\rightarrow\mathscr{B}(K)$ by $c\mapsto P_{K}\pi(c)P_{K}$ is onto, where $P_{K}$ is the projection onto $K$. This is equivalent to the statement that there is always an operator in $C$ sending a given orthonormal sequence to another given bounded sequence in $H$. Since the Calkin algebra is countably degree-1 saturated, it is reasonable to ask: Does every irreducible representation of the Calkin algebra have infinite transitivity? We produce a negative answer.
@@
Finally, we find an irreducible representation of the Calkin algebra which fails infinite transitivity, granting theorem 1. Compare this situation with that of [FA10], wherein the question of whether the commutant of $\mathscr{B}(H)$ in $\mathscr{B}(H)^{\mathscr{U}}$ is trivial also depends on the ultrafilter $\mathscr{U}$. In our case, for a diagonal pure state on $\mathscr{Q}(H)$ (or equivalently $\mathscr{B}(H)$), whether the GNS-representation admits infinite transitivity depends on the corresponding ultrafilter $\mathscr{U}$. Another recent instance of this ultrafilter-dependence phenomenon was in [ZA25].
@@ 1.4. Preliminaries
where $a_{0}\in C$, $a_{ijk}\in\tilde{C}$, $n\in\mathbb{N}$. A degree-$1$ condition of $C$ is an expression of the form
@@ 2.1. Quantum Filters
The beauty of maximal quantum filters is their natural bijective correspondence with pure states, i.e. $\varphi\mapsto\mathscr{F}_{\varphi}$ is a bijection from pure states to maximal quantum filters [FA19, 5.3.8]. However, the map $\varphi\mapsto\mathscr{F}_{\varphi}$ from states to quantum filters is almost never injective or surjective. For example, $\{1\}$ is always a quantum filter, but is only of the form $\mathscr{F}_{\varphi}$ when $A=\mathbb{C}$.
@@ 2.3. Equivalences
We state these two lemmas separately from theorem 2.8, as they apply to all states, not only pure ones. Also, lemma 2.7 does not require any saturation. Of particular interest to operator algebraists is the fact that these states have weak normality, a property first discovered for projections on the Calkin algebra by Bice in [BI13, 6.2]. That is, decreasing sequences $(a_{n})$ in $A_{+}$ have lower bounds $a\in A_{+}$ with $\varphi(a)=\lim_{n\rightarrow\infty}\varphi(a_{n})$.
@@
$\|(1-e_{\gamma})a-(1-e_{\gamma})\|$ $\leq\|1-e_{\gamma}\|\|a-a^{\prime}\|+\|e_{\gamma}(1-a^{\prime})-(1-a^{\prime})\|$
@@
so $(1-e_{\gamma})$ satisfies $t(x)$, so it is realized by $b\in\mathscr{F}_{\varphi}$ s.t. $b\ll a$ for all $a\in Y$. Since $Y$ was arbitrary, we are done.
@@
For $(4)\Leftrightarrow(7)$, we apply both directions of [BI13, 4.1].
@@
Without loss of generality, by the linearity of $\varphi$, whenever $\rho_{j}=\varphi$, assume $z_{j}=0$.
@@
$\rho^{\prime}_{j}=\begin{cases}\|\cdot\|,&\rho_{j}=\|\cdot\|;\\ \|a\cdot a\|,&\rho_{j}=\varphi;\end{cases}$
and fix the degree-$1$ norm type
@@
and notice $(\bar{b}_{m})$ satisfies $t^{\prime}(\bar{x})$. Since $C$ is countably degree-1 saturated, $t^{\prime}(\bar{x})$ is realized. When $\rho_{j}=\varphi$, we have
@@ 3. Infinite Transitivity
Let $t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\in K_{j}:j\in\mathbb{N}\}$ be a satisfiable countable type consisting of conditions above. To prove this, we will perform two conversions of $t(\bar{x})$.
@@
is a satisfiable generalized countable degree-1 norm type whose realizations realize $t^{\prime}(\bar{x})$, thus realize $t(\bar{x})$. To see this, we only need to check the case where $\rho_{j}=\|\cdot\|_{2}^{\varphi}$,
$\|P_{j}(\bar{b})\|_{2}^{\varphi}=\varphi(P_{j}(\bar{b})^{*}P_{j}(\bar{b}))^{\frac{1}{2}}\leq\|a_{\varphi}^{*}P_{j}(\bar{b})^{*}P_{j}(\bar{b})a_{\varphi}\|^{\frac{1}{2}}=\|P_{j}(\bar{b})a_{\varphi}\|\leq r_{j}.$
Therefore, by [FA19, 15.2.10], $t(\bar{x})$ is realized. ∎
@@
Let $K_{j}$ be a separable subspace of $H_{j}$ and enumerate an orthonormal basis $(\xi_{j})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\bigoplus_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{j})$. For each $j$, there exists some P-state $\varphi_{j}$ such that there is a lift $a_{j}+L_{\varphi_{j}}=\xi_{j}$. Similarly, there is some lift $b_{j}+L_{\varphi_{j}}=T\xi_{j}$. Define the type